### v2

# Bypassing the lattice BCS-BEC crossover in strongly correlated superconductors: resilient coherence from multiorbital physics

Abstract
Superconductivity emerges from the spatial coherence of a macroscopic condensate of Cooper pairs. Increasingly strong binding and localization of electrons into these pairs compromises the condensate’s phase stiffness, thereby limiting critical temperatures – a phenomenon known as the BCS-BEC crossover in lattice systems. In this study, we report on an enhancement of superconductivity beyond the limits of the lattice BCS-BEC crossover realized in a multiorbital model of alkali-doped fullerides (A3C60). We show how strong correlations and multiorbital effects lead into a localized superconducting regime characterized by a short coherence length but robust stiffness and a domeless rise in critical temperature with increasing pairing interaction. These insights are derived from the development of a theoretical framework to calculate the fundamental length scales of superconductors, namely the coherence length ($\xi_{0}$) and the London penetration depth ($\lambda_{\mathrm{L}}$), in microscopic theories and from first principles, even in presence of strong electron correlations.
coherence length, correlation length, penetration depth, fulleride, A3C60, alkali, doped, Buckyball, Buckminster, superconductivity, superconductor, fluctuations, s-wave, order parameter, depairing current, critical current, supercurrent, critical fields, finite-momentum pairing, helical phase, FFLO, Hund’s metal, strong coupling, multiorbital, Ginzburg-Landau, type-I, type-II, strong correlation, phase gradient, phase twist, stiffness, DMFT, Dynamical Mean-Field Theory, BCS to BEC crossover, BEC-BCS crossover, BEC, BCS, Uemura plot, superconducting dome
Introduction
The collective and phase coherent condensation of electrons into bound Cooper pairs leads to the emergence of superconductivity. This macroscopic coherence enables dissipationless charge currents, perfect diamagnetism, fluxoid quantization and technical applications (1, 2) ranging from electromagnets in particle accelerators to quantum computing hardware. Often, superconducting (SC) functionality is controlled by the critical surface spanned by critical magnetic fields, currents, and temperatures which a SC condensate can tolerate. Fundamentally, these are determined by the characteristic length scales of a superconductor – the London penetration depth, $\lambda_{\mathrm{L}}$, and the coherence length, $\xi_{0}$.
$\lambda_{\mathrm{L}}$ and $\xi_{0}$ quantify different aspects of the SC condensate: The penetration depth is the length associated with the mass term that the vector potential gains through the Anderson-Higgs mechanism (3). In consequence, magnetic fields decay exponentially over a distance $\lambda_{\mathrm{L}}$ inside a superconductor. Through this, $\lambda_{\mathrm{L}}$ is connected to the energy cost of order parameter (OP) phase variations and hence the SC stiffness $D_{\mathrm{s}}$. The coherence length, on the other hand, is the intrinsic length scale of OP amplitude variations and is associated with the amplitude Higgs mode. $\xi_{0}$ sets the scale below which amplitude and phase modes significantly couple such that spatial variations of the OP’s phase reduce its amplitude (3, 4).
In addition to influencing the macroscopic properties of superconductors, $\lambda_{\mathrm{L}}$ and $\xi_{0}$ play an important role to understand strongly correlated superconductors, as is epitomized in the Uemura plot (5, 6, 7, 8). For instance, the interplay of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ impacts critical temperatures (9), it is relevant for the pseudogap formation (10, 11, 12, 13), and it might underlie the light-enhancement of superconductivity (14, 15, 16, 17, 18). An important concept in this context is the BCS-BEC crossover phenomenology (19, 20, 21). It continuously connects the two limiting cases of weak-coupling Bardeen-Cooper-Schrieffer (BCS) superconductivity with weakly-bound and largely overlapping Cooper pairs to tightly-bound molecule-like pairs in the strong-coupling Bose-Einstein condensate (BEC) as the interaction strength or the density is varied (Fig. 1).
The BCS-BEC crossover has been studied in ultracold Fermi gases (20), low-density doped semiconductors (22), and is under debate for several unconventional superconductors (6, 8, 19, 23, 24, 25, 26). However, quasi-continuous systems, for instance Fermi gases, and strongly correlated superconducting solids show a crucially different behavior of how their SC properties, most importantly the critical temperature $T_{\mathrm{c}}$, change towards the strong coupling BEC limit. While $T_{\mathrm{c}}$ converges to a constant temperature TBEC for Fermi gases in a continuum (Fig. 1A), $T_{\mathrm{c}}$ can become arbitrarily small in strongly correlated lattice systems due to the quenching of kinetic energy (Fig. 1B). Since the movement of electron pairs, i.e., bosonic hopping, necessitates intermediate fermionic hopping, it becomes increasingly unfavorable for strong attractions. Thus, as Cooper pairs become localized on the scale of the lattice constant, the condensate’s stiffness and hence $T_{\mathrm{c}}$ are compromised (19, 21). Fig. 1 contrasts this generic BCS-BEC crossover picture for Fermi gases and correlated lattice systems in terms of the change of $T_{\mathrm{c}}$ and the pair size $\xi$ as a function of pairing strength. Due to the decrease of $T_{\mathrm{c}}$ in the BCS and BEC limits, a prominent dome-shape of $T_{\mathrm{c}}$ can be expected in the crossover region for solid materials. Because of this, recent experimental efforts to increase $T_{\mathrm{c}}$ concentrate on stabilizing SC materials in this region (19, 22, 24, 26).
Fig. 1: BCS-BEC crossover in Fermi gases vs. lattice systems. Evolution of the critical temperature $T_{\mathrm{c}}$ and Cooper pair size characterized by the coherence length $\xi$ in the BCS-BEC crossover (dark blue line) for (A) Fermi gases and (B) lattice systems. Both cases display a dome-shaped behavior of $T_{\mathrm{c}}$ in the intermediate crossover regime but behave qualitatively different in the strong coupling BEC phase: $T_{\mathrm{c}}$ remains finite in the Fermi gas (A) but approaches zero in the lattice case (B). During the crossover, $\xi$ is reduced to the order of interparticle spacing $r$ (A & B) and lattice constant $a$ (B). For the lattice system, we contrast the evolution towards the multiorbital strong coupling phase (light blue line) discussed in this article. Here, the localization of pairs differently affects the Bosonic hopping $t_{\mathrm{B}}$, as drawn in the insets on the right. In the lattice BEC limit, Bosonic hopping relies on a second-order process involving two fermionic hoppings, $t_{\mathrm{F}}$, and a virtual intermediate state with broken pairs that is inhibited by the strong attraction $g<0$. Consequently, this process and also $T_{\mathrm{c}}\propto t_{\mathrm{B}}=t_{\mathrm{F}}^{2}/|g|$ are quenched at large $|g|$ (19, 21). In the multiorbital strong coupling case, the local coexistence of paired and unpaired electrons fluctuating between different orbitals enables Bosonic hopping $t_{\mathrm{B}}$ as a first order process in $t_{\mathrm{F}}$ without any intermediate broken pair states. A second temperature scale $T^{*}$ is drawn in both panels as its splitting from $T_{\mathrm{c}}$ (corresponding to the opening of a pseudogap) marks the beginning of the crossover regime (19).
In this work, we demonstrate how multiorbital effects can enhance superconductivity beyond the expectations of the lattice BCS-BEC crossover phenomenology (as contrasted in Fig. 1B) with a model inspired by alkali-doped fullerides (A3C60 with A=K, Rb, Cs). The material family of A3C60 hosts exotic $s$-wave superconductivity of critical temperatures up to $T_{\mathrm{c}}=38$ K, being the highest temperatures among molecular superconductors (27, 28, 29, 30), and they possibly reach photo-induced SC at even higher temperatures (14, 16, 17, 31). In order to theoretically characterize the SC state, the knowledge of the intrinsic SC length scales is essential. While BCS theory and Eliashberg theory provide a microscopic description of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ for weakly correlated materials (32, 33, 34), their validity is unclear for superconductors with strong electron correlations. To the best of our knowledge, $\xi_{0}$ is generally not known from theory in strongly correlated materials. Only approaches to determine $\lambda_{\mathrm{L}}$ exist where an approximate, microscopic assessment of the SC stiffness from locally exact theories has been established, albeit neglecting vertex corrections (35, 36, 37, 38, 39).
Here, we introduce a novel theoretical framework to microscopically access $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from the tolerance of SC pairing to spatial OP variations. Central to our approach are calculations in the superconducting state under a constraint of finite-momentum pairing (FMP). Via Nambu-Gor’kov Green functions we get direct access to the superconducting OP and the depairing current $j_{\mathrm{dp}}$ which in turn yield $\xi_{0}$ and $\lambda_{\mathrm{L}}$. The FMP constraint is the SC analog to planar spin spirals applied to magnetic systems (40, 41, 42). As in magnetism, a generalized Bloch theorem holds that allows us to consider FMP without supercells; see section S2 in the supplementary text for a proof. As a result, our approach can be easily embedded in microscopic theories and ab initio approaches to tackle material-realistic calculations. In this work, we implement FMP in Dynamical Mean-Field theory (DMFT) (43) to treat strongly correlated superconductivity. In DMFT, the interacting many-body problem is solved by self-consistently mapping the lattice model onto a local impurity problem. By this, local correlations are treated exactly.
We apply the FMP-constrained in DMFT to A3C60 and we find $\xi_{0}$ and $\lambda_{\mathrm{L}}$ in line with experiment for model parameter ranges derived from ab initio estimates, validating our approach. For enhanced pairing interaction, we then reveal a multiorbital strong coupling SC state with minimal $\xi_{0}$ on the order of only 2 – 3 lattice constants, but with robust stiffness $D_{\mathrm{s}}$ and high $T_{\mathrm{c}}$ which increases with the pair interaction strength without a dome shape. This strong coupling SC state circumvents the lattice BCS-BEC crossover phenomenology. In the following, we motivate how the FMP constraint is linked to $\lambda_{\mathrm{L}}$ and $\xi_{0}$ with the subsequent application to the multiorbital model of A3C60.
Results
Superconductivity with the constraint of finite-momentum pairing
In most SC materials, Cooper pairs do not carry a finite center-of-mass momentum $\bm{q}=\bm{0}$. Yet, in presence of external fields, coexisting magnetism, or even spontaneously SC states with FMP, i.e., $\bm{q}\neq\bm{0}$, might arise (44, 45, 46, 47, 48) as originally conjectured in Fulde-Ferrel-Larkin-Ovchinnikov (FFLO) theory (49, 50, 51). Here, we enforce FMP states via constraints on the pair-center-of mass momentum $\bm{q}$ in order to access the characteristic SC length scales, $\xi_{0}$ and $\lambda_{\mathrm{L}}$, through an analysis of the momentum- and temperature-dependent OP and the closely connected charge supercurrent. We constrain the OP to be of the FMP form $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ corresponding to FF-type pairing (49), which is to be differentiated from pair density waves with amplitude modulations (50, 51, 52). We contrast the OP for zero and finite momentum in real and momentum space in the top panel of Fig. 2. For FMP, the OP’s phase is a helix winding around the direction of $\bm{q}$, while the OP for zero momentum pairing is simply a constant.
Before turning to the implementation in microscopic approaches, we motivate how the FMP constraint relates to $\lambda_{\mathrm{L}}$ and $\xi_{0}$. The Ginzburg-Landau (GL) framework provides an intuitive picture to this connection which we summarize here and discuss in detail in section S1 of the supplementary text. The GL low-order expansion of the free energy density $f_{\mathrm{GL}}$ in terms of the FMP-constrained OP reads
$f_{\mathrm{GL}}[\Psi_{\bm{q}}]=a|\Psi_{\bm{q}}|^{2}+\frac{b}{2}|\Psi_{\bm{q}}|^{4}+\frac{\hbar^{2}}{2m^{*}}q^{2}|\Psi_{\bm{q}}|^{2}$ (1)
with $q=|\bm{q}|$. $a$, $b$, and $m^{*}$ are the material and temperature dependent GL parameters. Here, the temperature dependent (GL) correlation length $\xi$ appears as the natural length scale of the amplitude mode ($\propto a$) and the kinetic energy term
$\xi(T)=\sqrt{\frac{\hbar^{2}}{2m^{*}|a|}}=\xi_{0}\left(1-\frac{T}{T_{\mathrm{c}}}\right)^{-\frac{1}{2}}$ (2)
with its zero-temperature value $\xi_{0}$ being the coherence length (4). The stationary point of Eq. (1) shows that the $\bm{q}$-dependent OP amplitude $\left|\Psi_{q}\right|=\left|\Psi_{0}\right|\sqrt{1-\xi^{2}q^{2}}$ decreases with increasing momentum $q$. For large enough $q$, SC order breaks down (i.e., $\Psi_{q}\to 0$) as the kinetic energy from phase modulations becomes comparable to the gain in energy from pairing. The length scale associated with this breakdown is $\xi$ and can, therefore, be inferred from the $q$-dependent OP suppression. We employ, here, the criterion $\xi=1/(\sqrt{2}|\bm{Q}|)$ with $\bm{Q}$ such that $|\Psi_{\bm{Q}}/\Psi_{0}|=1/\sqrt{2}$ (see section S5.A of the supplementary text for more information).
The finite center-of-mass momentum of the Cooper pairs entails a charge supercurrent $\bm{j}_{\bm{q}}\propto|\Psi_{\bm{q}}|^{2}\bm{q}$, c.f. Eq. (S18) in the supplementary text. This current density is a non-monotonous function of $q$ with a maximum called depairing current density $j_{\mathrm{dp}}$. It provides a theoretical upper bound to critical current densities, $j_{\mathrm{c}}$, measured in experiment. We note that careful design of SC samples is necessary for $j_{\mathrm{c}}$ reaching $j_{\mathrm{dp}}$ as its value crucially depends on sample geometry and defect densities (33, 53). $j_{\mathrm{dp}}$ is related to the London penetration depth via
$\lambda_{\mathrm{L}}(T)=\sqrt{\frac{\Phi_{0}}{3\sqrt{3}\pi\mu_{0}\xi j_{\mathrm{dp}}}}=\lambda_{\mathrm{L},0}\left(1-\left(\frac{T}{T_{\mathrm{c}}}\right)^{4}\right)^{-\frac{1}{2}}$ (3)
with the magnetic flux quantum $\Phi_{0}=h/2e$. The temperature dependence with the quartic power stated here is empirical (32, 33) and we find that it describes our calculations better than the linearized GL expectation as discussed in the supplementary text (section S5.B).
Fig. 2: Influence of finite-momentum pairing (FMP) constraint on the superconducting condensate. The top panel insets sketch the position and momentum space representation of the order parameter (OP) $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ in the zero-momentum (left, $q=0$) and finite-momentum pairing states (right, $q>0$). The main panels show (A) the momentum dependence of the OP modulus and (B) the supercurrent density $j_{q}=|\bm{j}_{\bm{q}}|$ as function of Cooper pair momentum $q=|\bm{q}|$ in reciprocal lattice units (r.l.u.). Gray lines indicate the points of extracting $\xi$ and $j_{\mathrm{dp}}$ (see text). The data shown are results for the A3C60 model (c.f. Eq. (7)) with interaction parameters $U/W=1.4$, $J/W=-0.04$ evaluated at different temperatures $T$ (color coded; see color bar with white triangular markers).
The GL analysis shows that the OP suppression and supercurrent induced by the FMP constraint connect to $\xi$ and $\lambda_{\mathrm{L}}$. In a microscopic description, we acquire the OP and supercurrent density from the Nambu-Gor’kov Green function $\mathcal{G}$
$\mathcal{G}_{\bm{q}}(\tau,\bm{k})$ $=-\langle T_{\tau}\psi_{\bm{k},\bm{q}}(\tau)\psi^{\dagger}_{\bm{k},\bm{q}}\rangle$
$=\begin{pmatrix}G_{\bm{q}}(\tau,\bm{k})&F_{\bm{q}}(\tau,\bm{k})\\ F^{*}_{\bm{q}}(\tau,\bm{k})&\bar{G}_{\bm{q}}(\tau,-\bm{k})\end{pmatrix}$ (4)
where $\psi_{\bm{k},\bm{q}}^{\dagger}=\left(c_{\bm{k}+\frac{\bm{q}}{2}\uparrow}^{\dagger}\ \ c_{-\bm{k}+\frac{\bm{q}}{2}\downarrow}\right)$ (orbital indices suppressed) are Nambu spinors that carry an additional dependence on $\bm{q}$ due to the FMP constraint. $G$ ($F$) denotes the normal (anomalous) Green function component for electrons ($G$) and holes ($\bar{G}$) in imaginary time $\tau$. For $s$-wave superconductivity as in the A3C60 model, we use the local anomalous Green function as the OP
$|\Psi_{\bm{q}}|\equiv[F^{\mathrm{loc}}_{\bm{q}}(\tau=0^{-})]_{\alpha\alpha}=\sum_{\bm{k}}\langle c_{\alpha\bm{k}+\frac{\bm{q}}{2}\uparrow}c_{\alpha-\bm{k}+\frac{\bm{q}}{2}\downarrow}\rangle\;,$ (5)
which is the same for all orbitals $\alpha$. The current density can be calculated via (c.f. Eq. (S37) in the supplementary text)
$\bm{j}_{\bm{q}}=\frac{2e}{N_{\bm{k}}}\sum_{\bm{k}}\mathrm{Tr}_{\alpha}\left[\underline{\bm{v}}(\bm{k})\underline{G}_{\bm{q}}\left(\tau=0^{-},\bm{k}-\frac{\bm{q}}{2}\right)\right]$ (6)
where $\hbar\underline{\bm{v}}=\nabla_{\bm{k}}\underline{h}(\bm{k})$ is the group velocity obtained from the one-electron Hamiltonian $\underline{h}(\bm{k})$ and the trace runs over the orbital indices of $\underline{\bm{v}}$ and $\underline{G}_{\bm{q}}$. Underlined quantities indicate matrices in orbital space, $N_{\bm{k}}$ is the number of momentum points and $e$ is the elementary charge. See Materials and Methods in the supplementary material for details on the DMFT-based implementation and sections S3, S4, and S5.B of the supplementary text for a derivation and discussion of Eq. (6).
The bottom panels of Fig. 2 show an example of our DMFT calculations which illustrates the $\bm{q}$-dependence of the OP amplitude and current density for different temperatures $T$. Throughout the paper, we choose $\bm{q}$ parallel to a reciprocal lattice vector $\bm{q}=q\bm{b}_{1}$. We find a monotonous suppression of the OP with increasing $q$. The supercurrent initially grows linearly with $q$, reaches its maximum $j_{\mathrm{dp}}$ and then collapses upon further increase of $q$. Thus, both $|\Psi_{q}|$ and $j_{q}$ behave qualitatively as expected from the GL description. For decreasing temperature, the point where the OP gets significantly suppressed moves towards larger momenta $q$ (smaller length scales $\xi$), while $j_{\mathrm{dp}}$ increases. We indicate the points where we extract $\xi$ and $j_{\mathrm{dp}}$ with gray circles connected by dashed lines.
Superconducting coherence in alkali-doped fullerides
We apply FMP superconductivity to study a degenerate three-orbital model $H=H_{\mathrm{kin}}+H_{\mathrm{int}}$, where $H_{\mathrm{kin}}$ is the kinetic energy and the electron-electron interaction is described by a local Kanamori-Hubbard interaction (54)
$H_{\mathrm{int}}=(U-3J)\frac{\hat{N}(\hat{N}-1)}{2}-J\left(2\hat{\bm{S}}^{2}+\frac{1}{2}\hat{\bm{L}}^{2}-\frac{5}{2}\hat{N}\right)$ (7)
with total number $\hat{N}$, spin $\hat{\bm{S}}$, and angular momentum operator $\hat{\bm{L}}$. The independent interactions are the intraorbital Hubbard term $U$ and Hund exchange $J$, as we use the rotational SU(2)$\times$×SO(3) symmetric parametrization. This model is often discussed in the context of Hund’s metal physics relevant to, e.g., transition-metal oxides like ruthenates with partially filled $t_{2g}$ shells (54, 55). In the special case of negative exchange energy $J<0$, it has been introduced to explain superconductivity in A3C60 materials (56, 57, 58). In fullerides, exotic $s$-wave superconductivity exists in proximity to a Mott-insulating (MI) (28, 29, 59) and a Jahn-Teller metallic phase (29, 60, 61, 62). The influence of strong correlation effects and inverted Hund’s coupling were shown to be essential for the SC pairing (56, 57, 58, 63) utilizing orbital fluctuations (64) in a Suhl-Kondo mechanism (60).
The inversion of $J$ seems unusual from the standpoint of atomic physics, where it dictates the filling of atomic shells via Hund’s rules. In A3C60, a negative $J$ is induced by the electronic system coupling to intramolecular Jahn-Teller phonon modes (63, 65, 66). As a result, Hund’s rules are inverted such that states which minimize first spin $\bm{S}$ and then angular momentum $\bm{L}$ are energetically most favorable, see the second term of Eq. (7) for $J<0$.
We connect the model to A3C60 by using an ab initio derived model of the half-filled $t_{1u}$ bands for $H_{\mathrm{kin}}$ with the bandwidth $W$ as the unit of energy ($W\approx 0.3$ – $0.5$ eV for Cs to K based A3C60) (58, 67). For the interaction, we take the first principles’ estimates of fixed $J/W=-0.04$ and varying $U/W$ (56, 57, 58, 61) to emulate unit cell volumes as resulting from the size of the different alkali dopants. We show the temperature dependence of $|\Psi_{q=0}|$, $\xi$, and $\lambda_{\mathrm{L}}$ for different $U/W$ in Fig. 3.
Fig. 3: Order parameter, correlation length, and penetration depth in A3C60. The temperature dependence of (A) the zero-momentum order parameter $|\Psi_{0}|$, (B) the correlation length $\xi$, and (C) the London penetration depth $\lambda_{\mathrm{L}}$ are shown for different ratios $U/W$. The data were obtained for fixed $J/W=-0.04$ as estimated from ab initio data. Fits to extract critical temperatures $T_{\mathrm{c}}$ and zero-temperature values $\xi_{0}$ and $\lambda_{\mathrm{L},0}$ are shown by dashed lines. The region of uncertainty to fit $T_{\mathrm{c}}$ is indicated by shaded regions.
[…]
Turning to the correlation length, we observe that, away from $T_{\mathrm{c}}$, $\xi$ is strongly reduced to only a few lattice constants ($a_{\mathrm{A}_{3}\mathrm{C}_{60}}\sim 14.2$ – 14.5 Å) by increasing $U$, i.e., pairing becomes very localized. At the same time, $\lambda_{\mathrm{L}}$ is enlarged. Hence, the condensate becomes much softer, i.e., there is a reduction of the SC stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon increasing $U$. Fitting Eqs. (2) and (3) to our data, we obtain zero-temperature values of $\xi_{0}=3$ – 10 nm and $\lambda_{\mathrm{L},0}=80$ – 120 nm. Comparing our results with experimental values of $\xi_{0}\sim 2$ – 4.5 nm and $\lambda_{\mathrm{L}}\sim 200$ – 800 nm (6, 7, 30, 68, 69), we see an almost quantitative match for $\xi_{0}$ and a qualitative match for $\lambda_{\mathrm{L}}$. Both experiment and theory consistently classify A3C60 as type II superconductors ($\lambda_{\mathrm{L}}\gg\xi_{0}$(69, 30).
We speculate that disorder and spontaneous orbital-symmetry breaking (62) in the vicinity of the Mott state could lead to a further reduction of $\xi_{0}$ as well as an increase of $\lambda_{\mathrm{L}}$ beyond what is found here for the pure system. This could bring our calculations with minimal $\xi_{0}=3$ nm closer to the experimental minimal coherence length of 2 nm revealed by measurements of large upper critical fields reaching up to a maximal $H_{\mathrm{c}2}=90$ T in Ref. 68 using $H_{\mathrm{c}2}=\Phi_{0}/(2\pi\xi_{0})$ with the flux quant $\Phi_{0}$.
Fig. 4: Superconducting state of the A3C60 model in the ($U$, $J$)-interaction space. (A) Critical temperature $T_{\mathrm{c}}$, zero temperature penetration depth $\lambda_{\mathrm{L},0}$, inverse square root of quasiparticle weight $Z$, coherence length $\xi_{0}$, and the statistical weight $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ of the local lowest energy states $\left|\upphi_{0}\right\rangle$ of the $N=2,3,4$ particle sectors obeying inverted Hund’s rules as a function of $U$ and $J$. Gray dots show original data points used for interpolation and the dashed line indicates a region where the proximity to the Mott state leads to a suppression of the superconducting state. There is no data point at the charge degeneracy line $U=2|J|$ in the lower left corner as marked by black ‘canceling’ lines. (B) Distribution of statistical weights $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ at four different interaction values $U$ and $J$. (See section S7 in the supplementary text for a listing of the eigenstates $\left|\upphi_{0}\right\rangle$ and their respective local eigenenergies $E_{n}^{N}$). Red (blue) bars denote the density matrix weight of ground states in the $N=3$ ($N=2,4$) particle sector, the sum of which is plotted in the last panel of A. (C) Exemplary depiction of representative lowest inverted Hund’s rule eigenstates. A delocalized doublon (electron pair) fluctuates between different orbitals due to correlated pair hopping $J$.
Circumvention of the lattice BCS-BEC crossover upon boosting inverted Hund’s coupling
The inverted Hund’s coupling is crucial for superconductivity in A3C60. This premise motivates us to explore in Fig. 4 the nature of the SC state in the interaction ($U$, $J$) phase space for $J<0$ beyond ab initio estimates.
[…]
The impact of $U$ on the SC state can be understood from the $U$-dependence of the London penetration depth: $\lambda_{\mathrm{L}}$ grows monotonously with $U$ and reaches its maximum close to the MI phase. Hence, the condensate is softest in the region where Mott physics is important and it becomes stiffer at smaller $U$. We find that this fits to the behavior of the effective bandwidth $W_{\mathrm{eff}}=ZW\propto D_{\mathrm{s}}$ where the quasiparticle weight $Z$ is suppressed upon approaching the Mott phase. The behavior of $Z$ shown in Fig. 4A confirms the qualitative connection $\lambda_{\mathrm{L}}\propto D_{\mathrm{s}}^{-1/2}\propto 1/\sqrt{Z}$ for $|J|>0.05$$W$. The $J$-dependence of $\lambda_{\mathrm{L}}$ is much weaker than the $U$-dependence, as can be seen in Fig. 4A and the corresponding line-cuts in Fig. 5.
$\xi_{0}$, in contrast, depends strongly on $J$. By just slightly increasing $|J|$ above the ab initio estimate of $|J|$, the SC state becomes strongly localized with a short coherence length on the order of 2 – 3 $a_{\mathrm{A}_{3}\mathrm{C}_{60}}$. Remarkably, the small value of $\xi_{0}$ is independent of $U$ and thus the proximity to the MI phase. The localization of the Cooper pairs with $\xi_{0}$ on the order of the lattice constant is reminiscent of a BEC-type SC state. However, the dome-shaped behavior, characteristic of the lattice BCS-BEC crossover, with decreasing $T_{\mathrm{c}}$ in the strong coupling limit is notably absent here. Instead, $T_{\mathrm{c}}$ still grows inside the plateau of minimal $\xi_{0}$ when increasing the effective pairing strength proportional to $|J|$ for fixed $U/W$ (c.f. Fig. 5A). Only by diagonally traversing the ($U$, $J$) phase space, it is possible to suppress $T_{\mathrm{c}}$ inside the short $\xi_{0}$ plateau similar to Ref. 61.
The reason for this circumvention of the lattice BCS-BEC phenomenology can be understood from an analysis of the local density matrix weights $\rho_{|\upphi_{n}\rangle}$, where $|\upphi_{n}\rangle$ refers to the eigenstates of the local Hamiltonian of our DMFT auxiliary impurity problem. We show $\rho_{|\upphi_{n}\rangle}$ of four different points in the interaction phase space in Fig. 4B. In the region of short $\xi_{0}$, the local density matrix is dominated by only eight states (red and blue bars) given by the “inverted Hund’s rule” ground states $|\upphi_{0}\rangle$ of the charge sectors with $N=2,3,4$ particles that are sketched in Fig. 4C. This can be seen in the last panel of Fig. 4A as the total weight of these eight states approaches one when entering the plateau of short $\xi_{0}$.
[…]
Correspondingly, as the superconducting state at $-J/W>0.05$ relies on direct transitions between the local inverted Hund’s ground states from filling $N=3$ to $N=2$ and $N=4$, the local Hubbard repulsion $U$ has to fulfill two requirements for superconductivity with appreciable critical temperatures. First, significant occupation of the $N=2$ and 4 states at half-filling requires that $U$ is not too large. Otherwise, the system turns Mott insulating (around $U/W\gtrsim 1.6$ for enhanced $|J|$) and the SC phase stiffness is reduced upon increasing $U$ towards the Mott limit. At the same time, a significant amount of statistical weight of the $N=3$ states demands that $U$ must not be too small either. We find that $\Delta_{\mathrm{at}}>0$ and thus $U>2|J|$ is necessary for robust SC pairing. At $\Delta_{\mathrm{at}}<0$ there is a predominance of $N=2$ and $4$ states, which couple kinetically only in second-order processes and which are susceptible to charge disproportionation (70). Our analysis shows that a sweet spot for a robust SC state with high phase stiffness and large $T_{\mathrm{c}}$ exists for $|J|$ approaching $U/2$.
Fig. 5: Crossover of superconducting properties between different interaction regimes in the A3C60 model. Critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, and stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L},0}^{-2}$ when approaching the multiorbital strong coupling state as function of $|J|/W$ at fixed $U/W=0.8$ (A) and when approaching the Mott insulating state as function of $U/W$ at fixed $J/W=-0.04$ (B). To indicate a general trend, each quantity is normalized to its maximal value within the line cut. The phenomenology of the SC regimes resembles the BCS to lattice BEC crossover in the latter case (B), which is distinct from the multiorbital strong coupling case (A) (c.f. summary listed in Table 1). Note that the $T_{\mathrm{c}}$ at $U/W=2$ in (B) corresponds to the transition from a Mott insulating state instead of a metallic phase, as visually differentiated by a dashed connecting line.
Discussion
We summarize the overall change from a weak-coupling BCS state to the multiorbital strong coupling SC state characterized by $T_{\mathrm{c}}$, $\xi_{0}$ and $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon enhancing $|J|$ in Fig. 5A and contrast it with the lattice BCS-BEC phenomenology in Table 1. Behavior as in the lattice BCS-BEC crossover can be found upon approaching the MI state (Fig. 5B), where strong local repulsion $U$ decreases $\xi_{0}$ and superconductivity compromises stiffness. Due to the MI state dominating over the SC state in our calculations, we cannot observe a proper $T_{\mathrm{c}}$ dome that is seen in the experimental phase diagram (29, 58). However, an additional analysis of the SC gap $\Delta$ and coupling strength in section S6 of the supplementary text indicates that our calculations are indeed in the vicinity of the crossover region. Thus, two distinct localized SC states exist in the multiorbital model A3C60— one facilitated by local Hubbard repulsion $U$ and the other by enhanced inverted Hund’s coupling $|J|$. One might speculate under which conditions THz driving or more generally photoexcitation (14, 16, 17, 31) could enhance $|J|$ and steer A3C60 into this high-$T_{\mathrm{c}}$ and short-$\xi_{0}$ strong coupling region, e.g., via quasiparticle trapping or displacesive meta-stability (71).
On general grounds, the bypassing of the usual lattice BCS-BEC scenario via multiorbital physics is promising for optimization of superconducting materials to achieve higher critical currents or temperatures. Generally, limits of accessible $T_{\mathrm{c}}$ are unknown with so far only a rigorous boundary existing for two-dimensional (2D) systems (72). An empirical upper bound to $T_{\mathrm{c}}$ emerges from the Uemura classification (5, 6, 7, 8) which compares $T_{\mathrm{c}}$ to the Fermi temperature $T_{\mathrm{F}}=E_{\mathrm{F}}/k_{\mathrm{B}}\propto D_{\mathrm{s}}$: the temperature $T_{\mathrm{BEC}}^{3\mathrm{D}}=T_{\mathrm{F}}/4.6$ of a three-dimensional (3D) non-interacting BEC. Most superconducting materials, including cuprate materials with the highest known $T_{\mathrm{c}}$ at ambient conditions, however, are not even close to this boundary; typically, $T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.1$ – $0.2$ for unconventional superconductors. Notable exceptions are monolayer FeSe ($T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.43$) (73), twisted graphitic systems ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.37$ – $0.57$) (23, 74), and 2D semiconductors at low carrier densities ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.36$ – $0.56$) (22) which all reach close to the 2D boundary $T_{\mathrm{c,lim}}^{2\mathrm{D}}/T_{\mathrm{BEC}}^{3\mathrm{D}}\approx 0.575$ (72). Only ultracold Fermi gases can be tuned very close to the optimal $T_{\mathrm{BEC}}$ (19, 20).
The empirical limitations of $T_{\mathrm{c}}$ in most unconventional superconductors with rather high densities make sense from the standpoint and constraints of the lattice BCS-BEC crossover. In this picture, high densities seem favorable for reaching high $T_{\mathrm{c}}$ as electrons can reach high intrinsic energy scales. Yet, lattice effects and their negative impact on $T_{\mathrm{c}}$ become also more pronounced, preventing $T_{\mathrm{c}}$ to reach $T_{\mathrm{BEC}}$. Possible routes to evade these constraints include quantum geometric and hybridization related band structure effects (75, 76, 77, 78, 79) as well as the multiorbital Hund’s interaction effects triggering localized stiffness uncovered here.
The framework to calculate the coherence length $\xi_{0}$ and the London penetration depth $\lambda_{\mathrm{L}}$ introduced, here, can be implemented to any Green function or density functional based approach to superconductivity without significant increase of the numerical complexity. Thus, our work opens the gate for “in silico” superconducting materials’ optimization targeting not only $T_{\mathrm{c}}$ but also $\xi_{0}$ and $\lambda_{\mathrm{L}}$. On a more fundamental level, availability of $\xi_{0}$ and $\lambda_{\mathrm{L}}$ rather than $T_{\mathrm{c}}$ alone can provide more constraints on possible pairing mechanisms through more rigorous theory-experiment comparisons, particularly in the domain of superconductors with strong electronic correlations.
Table 1: Characteristics of the BCS, lattice BEC, and multiorbital strong coupling limits. Behavior of the critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, superconducting stiffness $D_{\mathrm{s}}$, and the ratio of superconducting gap $\Delta$ to the Fermi energy $E_{\mathrm{F}}$ (coupling strength) in the BCS, lattice BEC limits (19) and “multiorbital strong coupling” limit. We indicate in the last row where we can find the respective behavior in the interaction plane of the A3C60 model. We do not observe the deep BEC limit for large $U/W$ and small $J/W$, but signatures indicating the onset of the BCS-BEC crossover (c.f. Figs. 4A and 5 as well as section S6 of the supplementary text).
  BCS limit Lattice BEC limit Multiorbital strong coupling
$\bm{T_{\mathrm{c}}}$ Increase with pairing interaction, $T_{\mathrm{c}}\propto\Delta(T=0)$ Decrease with pairing interaction, $T_{\mathrm{c}}\propto D_{\mathrm{s}}$ Increase with pairing interaction, $T_{\mathrm{c}}\propto\Delta,\,D_{\mathrm{s}}$
$\bm{\xi_{0}}$ Long ($\xi_{0}\ll a$) Short ($\xi_{0}\sim a)$ Short ($\xi_{0}\sim a)$
$\bm{D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}}$ Constant with pairing interaction Decreasing with pairing interaction Increasing/constant with pairing interaction
$\bm{\Delta}$/$\bm{E_{\mathrm{F}}}$ Small value ($\Delta/E_{\mathrm{F}}\ll 1$) Large value ($\Delta/E_{\mathrm{F}}\gtrsim 1$) Intermediate value ($\Delta/E_{\mathrm{F}}\sim\mathcal{O}(0.1)$)
Corresponding ($\bm{U}$, $\bm{J}$) region $\sim$ small $U/W$ $\sim$ small $|J|/W$ Onset at $\sim$ large $U/W$ $\sim$ small $|J|/W$ $\sim$ small to intermediate $U/W$ $\sim$ large $|J|/W$
Acknowledgements.
We thank M. Capone, A. Cavalleri, M. Eckstein, M. Katsnelson, Y. Iwasa, L. Mathey, T. Nomoto, G. Rai, S. Ryee, A. Toschi, and P. Werner for fruitful discussions. NW, AIL, and TW acknowledge funding by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) (EXC 2056, Project ID 390715994) and the DFG research unit FOR 5249 (‘QUAST’, Project No. 449872909). YN is grateful to funding via JSPS KAKENHI (No. JP23H04869 and JP21H01041). RA acknowledges funding by JSPS KAKENHI (No. JP19H05825). AIL is supported by the European Research Council via Synergy Grant 854843 - FASTCORR. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board under the project hhp00056 and provided on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure. Additional calculations were performed on the Physnet cluster at the University of Hamburg.