### v4

# Bypassing the lattice BCS–BEC crossover in strongly correlated superconductors through multiorbital physics

Abstract
Superconductivity emerges from the spatial coherence of a macroscopic condensate of Cooper pairs. Increasingly strong binding and localization of electrons into these pairs compromises the condensate’s phase stiffness, thereby limiting critical temperatures – a phenomenon known as the BCS–BEC crossover in lattice systems. In this study, we demonstrate enhanced superconductivity in a multiorbital model of alkali-doped fullerides (A3C60) that goes beyond the limits of the lattice BCS–BEC crossover. We identify that the interplay of strong correlations and multiorbital effects results in a localized superconducting state characterized by a short coherence length but robust stiffness and a domeless rise in critical temperature with increasing pairing interaction. To derive these insights, we introduce a new theoretical framework allowing us to calculate the fundamental length scales of superconductors, namely the coherence length ($\xi_{0}$) and the London penetration depth ($\lambda_{\mathrm{L}}$), even in presence of strong electron correlations.
coherence length, correlation length, penetration depth, fulleride, A3C60, alkali, doped, Buckyball, Buckminster, superconductivity, superconductor, fluctuations, s-wave, order parameter, depairing current, critical current, supercurrent, critical fields, finite-momentum pairing, helical phase, FFLO, Hund’s metal, strong coupling, multiorbital, Ginzburg–Landau, type-I, type-II, strong correlation, phase gradient, phase twist, stiffness, DMFT, Dynamical Mean-Field Theory, BCS to BEC crossover, BEC-BCS crossover, BEC, BCS, Uemura plot, superconducting dome
Introduction
The collective and phase coherent condensation of electrons into bound Cooper pairs leads to the emergence of superconductivity. This macroscopic coherence enables dissipationless charge currents, perfect diamagnetism, fluxoid quantization and technical applications 1, 2 ranging from electromagnets in particle accelerators to quantum computing hardware. Often, superconducting (SC) functionality is controlled by the critical surface spanned by critical magnetic fields, currents, and temperatures which a SC condensate can tolerate. Fundamentally, these are determined by the characteristic length scales of a superconductor – the London penetration depth, $\lambda_{\mathrm{L}}$, and the coherence length, $\xi_{0}$.
$\lambda_{\mathrm{L}}$ and $\xi_{0}$ quantify different aspects of the SC condensate: The penetration depth is the length associated with the mass term that the vector potential gains through the Anderson-Higgs mechanism 3. In consequence, magnetic fields decay exponentially over a distance $\lambda_{\mathrm{L}}$ inside a superconductor. Through this, $\lambda_{\mathrm{L}}$ is connected to the energy cost of order parameter (OP) phase variations and hence the SC stiffness $D_{\mathrm{s}}$. The coherence length, on the other hand, is the intrinsic length scale of OP amplitude variations and is associated with the amplitude Higgs mode. $\xi_{0}$ sets the scale below which amplitude and phase modes significantly couple such that spatial variations of the OP’s phase reduce its amplitude 3, 4.
In addition to influencing the macroscopic properties of superconductors, $\lambda_{\mathrm{L}}$ and $\xi_{0}$ play an important role to understand strongly correlated superconductors, as is epitomized in the Uemura plot 5, 6, 7, 8. For instance, the interplay of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ impacts critical temperatures 9, it is relevant for the pseudogap formation 10, 11, 12, 13, it influences magneto-thermal transport properties like the Nernst effect 8, 14, 15, and it might underlie the light-enhancement of superconductivity 15, 16, 17, 18, 19. An important concept in this context is the BCS–BEC crossover phenomenology 20, 21, 22, 23, 24. It continuously connects the two limiting cases of weak-coupling Bardeen–Cooper–Schrieffer (BCS) superconductivity with weakly-bound and largely overlapping Cooper pairs to tightly-bound molecule-like pairs in the strong-coupling Bose–Einstein condensate (BEC) as the interaction strength or the density is varied (Fig. 1).
The BCS–BEC crossover has been studied in ultracold Fermi gases 23, low-density doped semiconductors 25, and is under debate for several unconventional superconductors 6, 8, 24, 26, 27, 28, 29, 30. However, quasi-continuous systems, for instance Fermi gases, and strongly correlated superconducting solids show a crucially different behavior of how their SC properties, most importantly the critical temperature $T_{\mathrm{c}}$, change towards the strong coupling BEC limit. While $T_{\mathrm{c}}$ converges to a constant temperature $T_{\mathrm{BEC}}$ for Fermi gases in a continuum 31 (Fig. 1a), $T_{\mathrm{c}}$ can become arbitrarily small in strongly correlated lattice systems due to the quenching of kinetic energy (Fig. 1b). Since the movement of electron pairs, i.e., bosonic hopping, necessitates intermediate fermionic hopping, it becomes increasingly unfavorable for strong attractions. Thus, as Cooper pairs become localized on the scale of the lattice constant, the condensate’s stiffness and hence $T_{\mathrm{c}}$ are compromised 24, 21. Fig. 1 contrasts this generic BCS–BEC crossover picture for Fermi gases and correlated lattice systems in terms of the change of $T_{\mathrm{c}}$ and the pair size $\xi_{\mathrm{p}}$ as a function of pairing strength. Due to the decrease of $T_{\mathrm{c}}$ in the BCS and BEC limits, a prominent dome-shape of $T_{\mathrm{c}}$ can be expected in the crossover region for solid materials. Because of this, recent experimental efforts to increase $T_{\mathrm{c}}$ concentrate on stabilizing SC materials in this region 24, 25, 27, 30.
Fig. 1: BCS–BEC crossover in Fermi gases vs. lattice systems. Evolution of the critical temperature $T_{\mathrm{c}}$ and Cooper pair size $\xi_{\mathrm{p}}$ in the BCS–BEC crossover (dark blue line) for (a) Fermi gases and (b) lattice systems. Both cases display a dome-shaped behavior of $T_{\mathrm{c}}$ in the intermediate crossover regime but behave qualitatively different in the strong coupling BEC phase: $T_{\mathrm{c}}$ remains finite in the Fermi gas (a31 but approaches zero in the lattice case (b). During the crossover, $\xi_{\mathrm{p}}$ is reduced to the order of interparticle spacing $r$ (a & b) and lattice constant $a$ (b). For the lattice system, we contrast the evolution towards the multiorbital strong coupling phase (light blue line) discussed in this article. Here, the localization of pairs differently affects the bosonic hopping $t_{\mathrm{B}}$, as drawn in the insets on the right. In the lattice BEC limit, bosonic hopping relies on a second-order process involving two fermionic hoppings, $t_{\mathrm{F}}$, and a virtual intermediate state with broken pairs that is inhibited by the strong attraction $g<0$. Consequently, this process and also $T_{\mathrm{c}}\propto t_{\mathrm{B}}=t_{\mathrm{F}}^{2}/|g|$ are quenched at large $|g|$ 24, 21. In the multiorbital strong coupling case, the local coexistence of paired and unpaired electrons fluctuating between different orbitals enables bosonic hopping $t_{\mathrm{B}}$ as a first order process in $t_{\mathrm{F}}$ without any intermediate broken pair states. See Supplementary Note 7 for details on the involved states and energy diagram specific to the case of A3C60. A second temperature scale $T^{*}$ is drawn in both panels as its splitting from $T_{\mathrm{c}}$ (corresponding to the opening of a pseudogap) marks the beginning of the crossover regime 24.
In this work, we demonstrate how multiorbital effects32, 33 can enhance superconductivity beyond the expectations of the lattice BCS–BEC crossover phenomenology (as contrasted in Fig. 1b) with a model inspired by alkali-doped fullerides (A3C60 with A=K, Rb, Cs). The material family of A3C60 hosts exotic $s$-wave superconductivity of critical temperatures up to $T_{\mathrm{c}}=38$ K, being the highest temperatures among molecular superconductors 34, 35, 36, 37, and they possibly reach photo-induced SC at even higher temperatures 15, 17, 18, 38. In order to theoretically characterize the SC state, the knowledge of the intrinsic SC length scales is essential. While BCS theory and Eliashberg theory provide a microscopic description of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ for weakly correlated materials 39, 40, 41, their validity is unclear for superconductors with strong electron correlations. To the best of our knowledge, $\xi_{0}$ is generally not known from theory in strongly correlated materials. Only approaches to determine $\lambda_{\mathrm{L}}$ exist where an approximate, microscopic assessment of the SC stiffness from locally exact theories has been established, albeit neglecting vertex corrections 42, 43, 44, 45, 46, 47.
For this reason, we introduce a novel theoretical framework to microscopically access $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from the tolerance of SC pairing to spatial OP variations. Central to our approach are calculations in the superconducting state under a constraint of finite-momentum pairing (FMP). Via Nambu–Gor’kov Green functions we get direct access to the superconducting OP and the depairing current $j_{\mathrm{dp}}$ which in turn yield $\xi_{0}$ and $\lambda_{\mathrm{L}}$. The FMP constraint is the SC analog to planar spin spirals applied to magnetic systems 48, 49, 50. As in magnetism, a generalized Bloch theorem holds that allows us to consider FMP without supercells; see Supplementary Note 2 for a proof. As a result, our approach can be easily embedded in microscopic theories and ab initio approaches to tackle material-realistic calculations.
In this work, we implement FMP in Dynamical Mean-Field theory (DMFT) 51 to treat strongly correlated superconductivity. In DMFT, the interacting many-body problem is solved by self-consistently mapping the lattice model onto a local impurity problem. By this, local correlations are treated exactly. We apply the FMP-constrained DMFT to A3C60 where we find $\xi_{0}$ and $\lambda_{\mathrm{L}}$ in line with experiment for model parameter ranges derived from ab initio estimates, validating our approach. For enhanced pairing interaction, we then reveal a multiorbital strong coupling SC state with minimal $\xi_{0}$ on the order of only 2 – 3 lattice constants, but with robust stiffness $D_{\mathrm{s}}$ and high $T_{\mathrm{c}}$ which increases with the pair interaction strength without a dome shape. This strong coupling SC state is distinct to the lattice BCS–BEC crossover phenomenology showing promising routes to optimize superconductors with higher critical temperatures. We discuss this possibility in-depth after the presentation of our results.
The remaining paper is organized as follows: First, we motivate how the FMP constraint is linked to $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from phenomenological Ginzburg–Landau theory after which we summarize the microscopic approach from Green function methods. Technical details are available in the Supplementary Information as cross-referenced at relevant points. Subsequently, we discuss our results for the multiorbital model of A3C60. Readers primarily interested in the analysis of these results might skip the first part on the physical motivation for obtaining superconducting length scales from the FMP constraint.
Results
Extracting superconducting length scales from the constraint of finite-momentum pairing
In most SC materials, Cooper pairs do not carry a finite center-of-mass momentum $\bm{q}=\bm{0}$. Yet, in presence of external fields, coexisting magnetism, or even spontaneously SC states with FMP, i.e., $\bm{q}\neq\bm{0}$, might arise 52, 53, 54, 55, 56 as originally conjectured in Fulde–Ferrel–Larkin–Ovchinnikov (FFLO) theory 57, 58, 59. Here, we introduce a method to access the characteristic SC length scales $\xi_{0}$ and $\lambda_{\mathrm{L}}$ of strongly correlated materials through the calculation of FMP states with a manually constrained pair center-of-mass momentum $\bm{q}$40, 4, 60, 61, 62. For this, we enforce the OP to be of the FMP form $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ corresponding to FF-type pairing 57, which is to be differentiated from pair density waves with amplitude modulations 58, 59, 63. We contrast the OP for zero and finite momentum in real and momentum space in the top panel of Fig. 2. For FMP, the OP’s phase is a helix winding around the direction of $\bm{q}$, while the OP for zero-momentum pairing is simply a constant.
Before turning to the implementation in microscopic approaches, we motivate how the FMP constraint relates to $\lambda_{\mathrm{L}}$ and $\xi_{0}$. The Ginzburg–Landau (GL) framework provides an intuitive picture to this connection40, 4, 61 which we summarize here and discuss in detail in Supplementary Note 1, including a derivation of the following equations. The GL low-order expansion of the free energy density $f_{\mathrm{GL}}$ in terms of the FMP-constrained OP reads
$f_{\mathrm{GL}}[\Psi_{\bm{q}}]=\alpha|\Psi_{\bm{q}}|^{2}+\frac{b}{2}|\Psi_{\bm{q}}|^{4}+\frac{\hbar^{2}}{2m^{*}}q^{2}|\Psi_{\bm{q}}|^{2}$ (1)
with $q=|\bm{q}|$. $\alpha$, $b$, and $m^{*}$ are the material and temperature dependent GL parameters. Here, the temperature-dependent (GL) correlation length $\xi(T)$ appears as the natural length scale of the amplitude mode ($\propto|\alpha|$) and the kinetic energy term
$\xi(T)=\sqrt{\frac{\hbar^{2}}{2m^{*}|\alpha|}}=\xi_{0}\left(1-\frac{T}{T_{\mathrm{c}}}\right)^{-\frac{1}{2}}$ (2)
with its zero-temperature value $\xi_{0}$ being the coherence length 4. The stationary point of Eq. (1) shows that the $\bm{q}$-dependent OP amplitude $\left|\Psi_{\bm{q}}\right|=\left|\Psi_{0}\right|\sqrt{1-\xi^{2}q^{2}}$ ($T$-dependence suppressed) decreases with increasing momentum $q$. For large enough $q$, SC order breaks down (i.e., $\Psi_{\bm{q}}\to 0$) as the kinetic energy from phase modulations becomes comparable to the gain in energy from pairing. The length scale associated with this breakdown is $\xi(T)$ and can, therefore, be inferred from the $q$-dependent OP suppression. We employ, here, the criterion $\xi(T)=1/(\sqrt{2}|\bm{Q}|)$ with $\bm{Q}$ such that $|\Psi_{\bm{Q}}(T)/\Psi_{0}(T)|=1/\sqrt{2}$ for fixed $T$ (see Supplementary Note 5-A for more information).
The finite center-of-mass momentum of the Cooper pairs entails a charge supercurrent $\bm{j}_{\bm{q}}\propto|\Psi_{\bm{q}}|^{2}\bm{q}$, c.f. Eq. (8) in Supplementary Note 1. This current density is a non-monotonous function of $q$ with a maximum called depairing current density $j_{\mathrm{dp}}$. It provides a theoretical upper bound to critical current densities, $j_{\mathrm{c}}$, measured in experiment. We note that careful design of SC samples is necessary for $j_{\mathrm{c}}$ reaching $j_{\mathrm{dp}}$ as its value crucially depends on sample geometry and defect densities 40, 61, 64. As the current is related to the London penetration depth $\lambda_{\mathrm{L}}$ through the second London equation $\mu_{0}\bm{j}=-\lambda_{\mathrm{L}}^{-2}\bm{A}$, we can determine $\lambda_{\mathrm{L}}$ from $j_{\mathrm{dp}}$ in GL theory via
$\lambda_{\mathrm{L}}(T)=\sqrt{\frac{\Phi_{0}}{3\sqrt{3}\pi\mu_{0}\xi(T)j_{\mathrm{dp}}(T)}}=\lambda_{\mathrm{L},0}\left(1-\left(\frac{T}{T_{\mathrm{c}}}\right)^{4}\right)^{-\frac{1}{2}}$ (3)
with the magnetic flux quantum $\Phi_{0}=h/2e$. The temperature dependence with the quartic power stated here is empirical 39, 40 and we find that it describes our calculations better than the linearized GL expectation as discussed in Supplementary Note 5-B. Note that taking the $q\to 0$ limit in our approach is related to linear-response-based methods to calculate $\lambda_{\mathrm{L}}$ or, equivalently, the stiffness $D_{\mathrm{s}}$ 62, 47.
Fig. 2: Influence of finite-momentum pairing (FMP) constraint on the superconducting condensate. The top panel insets sketch the position and momentum space representation of the order parameter (OP) $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ in the zero-momentum (left, $q=0$) and finite-momentum pairing states (right, $q>0$). The main panels show (a) the momentum dependence of the OP modulus and (b) the supercurrent density $j_{q}=|\bm{j}_{\bm{q}}|$ as function of Cooper pair momentum $q=|\bm{q}|$ in reciprocal lattice units (r.l.u.). Gray lines indicate the points of extracting $\xi$ and $j_{\mathrm{dp}}$ (see text). The data shown are results for the A3C60 model (c.f. Eq. (7)) with lattice constant $a$ and interaction parameters $U/W=1.4$, $J/W=-0.04$ evaluated at different temperatures $T$ (color coded; see color bar with white triangular markers).
The GL analysis shows that the OP suppression and supercurrent induced by the FMP constraint connect to $\xi_{0}$ and $\lambda_{\mathrm{L}}$. In a microscopic description, we acquire the OP and supercurrent density from the Nambu–Gor’kov Green function
$\mathcal{G}_{\bm{q}}(\tau,\bm{k})$ $=-\langle T_{\tau}\psi_{\bm{k},\bm{q}}(\tau)\psi^{\dagger}_{\bm{k},\bm{q}}\rangle$
$=\begin{pmatrix}G_{\bm{q}}(\tau,\bm{k})&F_{\bm{q}}(\tau,\bm{k})\\ F^{*}_{\bm{q}}(\tau,\bm{k})&\bar{G}_{\bm{q}}(\tau,-\bm{k})\end{pmatrix}$ (4)
where $\psi_{\bm{k},\bm{q}}^{\dagger}=\left(c_{\bm{k}+\frac{\bm{q}}{2}\uparrow}^{\dagger}\ \ c_{-\bm{k}+\frac{\bm{q}}{2}\downarrow}\right)$ (orbital indices suppressed) are Nambu spinors that carry an additional dependence on $\bm{q}$ due to the FMP constraint. $G$ ($F$) denotes the normal (anomalous) Green function component for electrons ($G$) and holes ($\bar{G}$) in imaginary time $\tau$. For $s$-wave superconductivity as in A3C60 65, 66, 67, 68, we use the orbital-diagonal, local anomalous Green function as the OP
$|\Psi_{\bm{q}}|\equiv[F^{\mathrm{loc}}_{\bm{q}}(\tau=0^{-})]_{\alpha\alpha}=\sum_{\bm{k}}\langle c_{\alpha\bm{k}+\frac{\bm{q}}{2}\uparrow}c_{\alpha-\bm{k}+\frac{\bm{q}}{2}\downarrow}\rangle\;,$ (5)
which is the same for all orbitals $\alpha$. This allows us to work with a single-component OP. The current density can be calculated via (c.f. Eq. (26) in the Supplementary Information)
$\bm{j}_{\bm{q}}=\frac{2e}{N_{\bm{k}}}\sum_{\bm{k}}\mathrm{Tr}_{\alpha}\left[\underline{\bm{v}}(\bm{k})\underline{G}_{\bm{q}}\left(\tau=0^{-},\bm{k}-\frac{\bm{q}}{2}\right)\right]$ (6)
where $\hbar\underline{\bm{v}}=\nabla_{\bm{k}}\underline{h}(\bm{k})$ is the group velocity obtained from the one-electron Hamiltonian $\underline{h}(\bm{k})$ and the trace runs over the orbital indices of $\underline{\bm{v}}$ and $\underline{G}_{\bm{q}}$. Underlined quantities indicate matrices in orbital space, $N_{\bm{k}}$ is the number of momentum points, and $e$ is the elementary charge. See Methods and Supplementary Note 4 for details on the DMFT-based implementation and Supplementary Notes 3 for a derivation and discussion of Eq. (6).
The bottom panels of Fig. 2 show an example of our DMFT calculations which illustrates the $\bm{q}$-dependence of the OP amplitude and current density for different temperatures $T$. Throughout the paper, we choose $\bm{q}$ parallel to a reciprocal lattice vector $\bm{q}=q\bm{b}_{1}$. We find a monotonous suppression of the OP with increasing $q$. The supercurrent initially grows linearly with $q$, reaches its maximum $j_{\mathrm{dp}}$ and then collapses upon further increase of $q$. Thus, both $|\Psi_{q}|$ and $j_{q}$ behave qualitatively as expected from the GL description. For decreasing temperature, the point where the OP gets significantly suppressed moves towards larger momenta $q$ (smaller length scales $\xi(T)$), while $j_{\mathrm{dp}}(T)$ increases. We indicate the points where we extract $\xi(T)$ and $j_{\mathrm{dp}}(T)$ with gray circles connected by dashed lines.
Superconducting coherence in alkali-doped fullerides
We apply FMP superconductivity to study a degenerate three-orbital model $H=H_{\mathrm{kin}}+H_{\mathrm{int}}$, where $H_{\mathrm{kin}}$ is the kinetic energy and the electron-electron interaction is described by a local Kanamori–Hubbard interaction 69
$H_{\mathrm{int}}=(U-3J)\frac{\hat{N}(\hat{N}-1)}{2}-J\left(2\hat{\bm{S}}^{2}+\frac{1}{2}\hat{\bm{L}}^{2}-\frac{5}{2}\hat{N}\right)$ (7)
with total number $\hat{N}$, spin $\hat{\bm{S}}$, and angular momentum operator $\hat{\bm{L}}$. The independent interactions are the intraorbital Hubbard term $U$ and Hund exchange $J$, as we use the SU(2)$\times$×SO(3) symmetric parametrization. This model is often discussed in the context of Hund’s metal physics relevant to, e.g., transition-metal oxides like ruthenates with partially filled $t_{2g}$ shells 69, 70. In the special case of negative exchange energy $J<0$, it has been introduced to explain superconductivity in A3C60 materials 65, 66, 67, 68. In fullerides, exotic $s$-wave superconductivity exists in proximity to a Mott-insulating (MI) 35, 36, 71 and a Jahn–Teller metallic phase 36, 72, 73, 74. The influence of strong correlation effects and inverted Hund’s coupling were shown to be essential for the SC pairing 65, 66, 67, 68, 75 utilizing orbital fluctuations 76 in a Suhl–Kondo mechanism 72. In the weakly correlated regime of $U\to 0$ and small $J$, the system follows BCS phenomenology 66.
The inversion of $J$ seems unusual from the standpoint of atomic physics, where it dictates the filling of atomic shells via Hund’s rules. In A3C60, a negative $J$ is induced by the electronic system coupling to intramolecular Jahn–Teller phonon modes 75, 77, 78, 68. As a result, Hund’s rules are inverted such that states which minimize first spin $\bm{S}$ and then angular momentum $\bm{L}$ are energetically most favorable, see the second term of Eq. (7) for $J<0$.
We connect the model to A3C60 by using an ab initio derived model for the kinetic energy 79
$\begin{split}H_{\mathrm{kin}}&=\sum_{ij}\sum_{\alpha\gamma\sigma}t_{\alpha\gamma}(\bm{R}_{ij})c^{\dagger}_{i\alpha\sigma}c_{j\gamma\sigma}\end{split}$ (8)
where $t_{\alpha\gamma}(\bm{R}_{ij})$ is the hopping amplitude between half-filled $t_{1u}$ orbitals ($\alpha,\gamma=1,2,3$) of C60 molecules on sites connected by lattice vector $\bm{R}_{ij}$. We take the bandwidth $W$ as the unit of energy ($W\approx 0.3$ – $0.5$ eV for Cs to K based A3C6068, 79, see Supplemental Note 4-D for further details. For the interaction, we take the first principles’ estimates of fixed $J/W=-0.04$ and varying $U/W$ 65, 66, 67, 68, 73 to emulate unit cell volumes as resulting from the size of different alkali dopants. We solve this Hamiltonian using DMFT, which explicitly takes into account superconducting order. Through the momentum dependence of $|\Psi_{\bm{q}}(T)|$ and ${\bm{j}}_{\bm{q}}(T)$, $\xi(T)$ and $\lambda_{\mathrm{L}}(T)$ can be extracted as discussed in the previous subsection. We show the derived temperature dependence of $|\Psi_{q=0}(T)|$, $\xi(T)$, and $\lambda_{\mathrm{L}}(T)$ for different $U/W$ in Fig. 3.
Fig. 3: Order parameter, correlation length, and penetration depth in A3C60. The temperature dependence of (a) the zero-momentum order parameter $|\Psi_{0}|$, (b) the correlation length $\xi$, and (c) the London penetration depth $\lambda_{\mathrm{L}}$ are shown for different ratios $U/W$. The data were obtained for fixed $J/W=-0.04$ as estimated from ab initio data. Length scales are given in units of the lattice constant $a$. Fits to extract critical temperatures $T_{\mathrm{c}}$ and zero-temperature values $\xi_{0}$ and $\lambda_{\mathrm{L},0}$ are shown by dashed lines. The region of uncertainty to fit $T_{\mathrm{c}}$ is indicated by shaded regions.
[…]
Turning to the correlation length, we observe that, away from $T_{\mathrm{c}}$, $\xi(T)$ is strongly reduced to only a few lattice constants ($a\sim 14.2$ – 14.5 Å) by increasing $U$, i.e., pairing becomes very localized. At the same time, $\lambda_{\mathrm{L}}(T)$ is enlarged. Hence, the condensate becomes much softer as there is a reduction of the SC stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon increasing $U$. Fitting Eqs. (2) and (3) to our data, we obtain zero-temperature values of $\xi_{0}=3$ – 10 nm and $\lambda_{\mathrm{L},0}=80$ – 120 nm. Comparing our results with experimental values of $\xi_{0}\sim 2$ – 4.5 nm and $\lambda_{\mathrm{L}}\sim 200$ – 800 nm 6, 7, 37, 80, 81, we see an almost quantitative match for $\xi_{0}$ and a qualitative match for $\lambda_{\mathrm{L}}$. Both experiment and theory consistently classify A3C60 as type II superconductors ($\lambda_{\mathrm{L}}\gg\xi_{0}$81, 37.
We speculate that disorder and spontaneous orbital-symmetry breaking 74 in the vicinity of the Mott state could lead to a further reduction of $\xi_{0}$ as well as an increase of $\lambda_{\mathrm{L}}$ beyond what is found here for the pure system. This could bring our calculations with minimal $\xi_{0}=3$ nm closer to the experimental minimal coherence length of 2 nm revealed by measurements of large upper critical fields reaching up to a maximal $H_{\mathrm{c}2}=90$ T in Ref. 80 using $H_{\mathrm{c}2}=\Phi_{0}/(2\pi\xi_{0})$ with the flux quant $\Phi_{0}$.
Fig. 4: Superconducting state of the A3C60 model in the ($U$, $J$)-interaction space. (a) Critical temperature $T_{\mathrm{c}}$, zero temperature penetration depth $\lambda_{\mathrm{L},0}$, inverse square root of quasiparticle weight $Z$, coherence length $\xi_{0}$, and the statistical weight $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ of the local lowest energy states $\left|\upphi_{0}\right\rangle$ of the $N=2,3,4$ particle sectors obeying inverted Hund’s rules as a function of $U$ and $J$. Gray dots show original data points used for interpolation and the dashed line indicates a region where the proximity to the Mott state leads to a suppression of the superconducting state. There is no data point at the charge degeneracy line $U=2|J|$ in the lower left corner as marked by black ‘canceling’ lines. (b) Distribution of statistical weights $\rho_{\left|\upphi_{n}\right\rangle}^{N}$ at four different interaction values $U$ and $J$. (See Supplementary Note 7 for a listing of the eigenstates $\left|\upphi_{n}\right\rangle$ and their respective local eigenenergies $E_{n}^{N}$). Red (blue) bars denote the density matrix weight of ground states in the $N=3$ ($N=2,4$) particle sector, the sum of which is plotted in the last panel of a. (c) Exemplary depiction of representative lowest inverted Hund’s rule eigenstates. A delocalized doublon (electron pair) fluctuates between different orbitals due to correlated pair hopping $J$.
Circumvention of the lattice BCS–BEC crossover upon boosting inverted Hund’s coupling
The inverted Hund’s coupling is crucial for superconductivity in A3C60. This premise motivates us to explore in Fig. 4 the nature of the SC state in the interaction ($U$, $J$) phase space for $J<0$ beyond ab initio estimates.
[…]
The impact of $U$ on the SC state can be understood from the $U$-dependence of the London penetration depth: $\lambda_{\mathrm{L}}$ grows monotonously with $U$ and reaches its maximum close to the MI phase. Hence, the condensate is softest in the region where Mott physics is important and it becomes stiffer at smaller $U$. We find that this fits to the behavior of the effective bandwidth $W_{\mathrm{eff}}=ZW\propto D_{\mathrm{s}}$ where the quasiparticle weight $Z$ is suppressed upon approaching the Mott phase. The behavior of $Z$ shown in Fig. 4a confirms the qualitative connection $\lambda_{\mathrm{L}}\propto D_{\mathrm{s}}^{-1/2}\propto 1/\sqrt{Z}$ for $|J|>0.05$$W$. The $J$-dependence of $\lambda_{\mathrm{L}}$ is much weaker than the $U$-dependence, as can be seen in Fig. 4a and the corresponding line-cuts in Fig. 5.
$\xi_{0}$, in contrast, depends strongly on $J$. By just slightly increasing $|J|$ above the ab initio estimate of $|J|$, the SC state becomes strongly localized with a short coherence length on the order of 2 – 3 $a$. Remarkably, the small value of $\xi_{0}$ is independent of $U$ and thus the proximity to the MI phase. The localization of the condensate with $\xi_{0}$ on the order of the lattice constant is reminiscent of a crossover to the BEC-type SC state. However, the dome-shaped behavior, characteristic of the lattice BCS–BEC crossover 27, 29, with decreasing $T_{\mathrm{c}}$ in the strong coupling limit is notably absent here. Instead, $T_{\mathrm{c}}$ still grows inside the plateau of minimal $\xi_{0}$ when increasing the effective pairing strength proportional to $|J|$ for fixed $U/W$ (c.f. Fig. 5a). Only by diagonally traversing the ($U$, $J$) phase space, it is possible to suppress $T_{\mathrm{c}}$ inside the short $\xi_{0}$ plateau with a dome structure as shown, e.g., in Ref. 73.
The reason for this circumvention of the lattice BCS–BEC phenomenology can be understood from an analysis of the local density matrix weights $\rho_{|\upphi_{n}\rangle}$, where $|\upphi_{n}\rangle$ refers to the eigenstates of the local Hamiltonian of our DMFT auxiliary impurity problem. We show $\rho_{|\upphi_{n}\rangle}$ of four different points in the interaction phase space in Fig. 4b. In the region of short $\xi_{0}$, the local density matrix is dominated by only eight states (red and blue bars) given by the “inverted Hund’s rule” ground states $|\upphi_{0}\rangle$ of the charge sectors with $N=2,3,4$ particles that are sketched in Fig. 4c. This can be seen in the last panel of Fig. 4a as the total weight of these eight states approaches one when entering the plateau of short $\xi_{0}$.
By increasing $|J|$, the system is driven into a strong coupling phase where local singlets are formed as Cooper pair precursors 77 while electronic hopping is not inhibited. On the contrary, hopping between the $N=2,3,4$ ground states is even facilitated via large negative $J$. It does so by affecting two different energy scales (c.f. Supplementary Note 7): Enhancing $J<0$ reduces the atomic gap $\Delta_{\mathrm{at}}=E_{0}^{N=4}+E_{0}^{N=2}-2E_{0}^{N=3}=U-2|J|$ 69 relevant to charge excitations and thereby supports hopping. A higher negative Hund’s exchange simultaneously increases the energy $\Delta E=2|J|$ necessary to break up the orbital singlets within a fixed charge sector. As a result, unpaired electrons in the $N=3$ state become more itinerant while the local Cooper pair binding strength increases. Since the hopping is not reduced, the SC stiffness is not compromised by larger $|J|$. This two-faced role or Janus effect of negative Hund’s exchange, that localizes Cooper pairs but delocalizes electrons, can be understood as a competition of a Mott and a charge disproportionated insulator giving way for a mixed-valence metallic state in between 82.
Correspondingly, as the superconducting state at $-J/W>0.05$ relies on direct transitions between the local inverted Hund’s ground states from filling $N=3$ to $N=2$ and $N=4$, the local Hubbard repulsion $U$ has to fulfill two requirements for superconductivity with appreciable critical temperatures. First, significant occupation of the $N=2$ and 4 states at half-filling requires that $U$ is not too large. Otherwise, the system turns Mott insulating (around $U/W\gtrsim 1.6$ for enhanced $|J|$) and the SC phase stiffness is reduced upon increasing $U$ towards the Mott limit. At the same time, a significant amount of statistical weight of the $N=3$ states demands that $U$ must not be too small either. We find that $\Delta_{\mathrm{at}}>0$ and thus $U>2|J|$ is necessary for robust SC pairing. At $\Delta_{\mathrm{at}}<0$ there is a predominance of $N=2$ and $4$ states, which couple kinetically only in second-order processes and which are susceptible to charge disproportionation 82. Our analysis shows that a sweet spot for a robust SC state with high phase stiffness and large $T_{\mathrm{c}}$ exists for $|J|$ approaching $U/2$.
Fig. 5: Crossover of superconducting properties between different interaction regimes in the A3C60 model. Critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, and stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L},0}^{-2}$ when approaching the multiorbital strong coupling state as function of $|J|/W$ at fixed $U/W=0.8$ (a) and when approaching the Mott insulating state as function of $U/W$ at fixed $J/W=-0.04$ (b). To indicate a general trend, each quantity is normalized to its maximal value within the line cut. The phenomenology of the SC regimes resembles the BCS to lattice BEC crossover when approaching the Mott insulator (b), which is distinct from the multiorbital strong coupling case (a) (c.f. summary listed in Table 1). Note that the $T_{\mathrm{c}}$ at $U/W=2$ in (b) corresponds to the transition from a Mott insulating state instead of a metallic phase, as visually differentiated by a dashed connecting line.
Discussion
We summarize the overall change from a weak-coupling BCS state to the multiorbital strong coupling SC state characterized by $T_{\mathrm{c}}$, $\xi_{0}$, and $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon enhancing $|J|$ in Fig. 5a and contrast it with the lattice BCS–BEC phenomenology in Table 1. Behavior as in the lattice BCS–BEC crossover can be found upon approaching the MI state (Fig. 5b), where strong local repulsion $U$ decreases $\xi_{0}$ and superconductivity compromises stiffness. As discussed above, we cannot observe a proper $T_{\mathrm{c}}$ dome that is seen in the experimental phase diagram due to the MI state dominating over the SC state in our calculations for fixed $J/W=-0.04$ 36, 67, 68. However, an additional analysis of the SC gap $\Delta$ and coupling strength in Supplementary Note 6 indicates that our calculations are indeed in the vicinity of the crossover region. Diagonally traversing the ($U$,$J$) interaction plane by increasing $U$ for $J=-\gamma U$ ($\gamma\leq 0.5$) leads to a similar conclusion, although a $T_{\mathrm{c}}$ dome can emerge more easily 73 as charge excitations controlled by $\Delta_{\mathrm{at}}$ are suppressed at a slower pace.
Thus, two distinct localized SC states exist in the multiorbital model A3C60– one facilitated by local Hubbard repulsion $U$ and the other by enhanced inverted Hund’s coupling $|J|$. One might speculate under which conditions THz driving or more generally photoexcitation 15, 17, 18, 38 could enhance $|J|$ and steer A3C60 into this high-$T_{\mathrm{c}}$ and short-$\xi_{0}$ strong coupling region, e.g., via quasiparticle trapping or displacesive meta-stability 83. Further experimental characterization via observables susceptible to changes in $\xi_{0}$ and $\lambda_{\mathrm{L}}$, like critical fields, currents (see Supplementary Note 1-B), or thermoelectric and thermal transport coefficients 14, 8, can lash down the possibility of this scenario. Twisted bilayer graphene might be another platform to host the mechanism proposed in this work as an inverted Hund’s pairing similar to that in A3C60 is currently under discussion 84, 85, 86.
On general grounds, the bypassing of the usual lattice BCS–BEC scenario via multiorbital physics is promising for optimization of superconducting materials to achieve higher critical currents or temperatures. Generally, limits of accessible $T_{\mathrm{c}}$ are unknown with so far only a rigorous boundary existing for two-dimensional (2D) systems 87, 88. An empirical upper bound to $T_{\mathrm{c}}$ emerges from the Uemura classification 5, 6, 7, 8 which compares $T_{\mathrm{c}}$ to the Fermi temperature $T_{\mathrm{F}}=E_{\mathrm{F}}/k_{\mathrm{B}}\propto D_{\mathrm{s}}$: the temperature $T_{\mathrm{BEC}}^{3\mathrm{D}}=T_{\mathrm{F}}/4.6$ of a three-dimensional (3D) non-interacting BEC. Most superconducting materials, including cuprate materials with the highest known $T_{\mathrm{c}}$ at ambient conditions, however, are not even close to this boundary; typically, $T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.1$ – $0.2$ for unconventional superconductors. Notable exceptions are monolayer FeSe89 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.43$), twisted graphitic systems26, 84 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.37$ – $0.57$), and 2D semiconductors at low carrier densities25 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.36$ – $0.56$) which all reach close to the 2D boundary $T_{\mathrm{c,lim}}^{2\mathrm{D}}/T_{\mathrm{BEC}}^{3\mathrm{D}}\approx 0.575$ 87. Only ultracold Fermi gases can be tuned very close to the optimal $T_{\mathrm{BEC}}$ 24, 23.
The empirical limitations of $T_{\mathrm{c}}$ in most unconventional superconductors with rather high densities make sense from the standpoint and constraints of the lattice BCS–BEC crossover. In this picture, high densities seem favorable for reaching high $T_{\mathrm{c}}$ as electrons can reach high intrinsic energy scales. Yet, lattice effects and their negative impact on the kinetic energy of superconducting carriers become also more pronounced, preventing $T_{\mathrm{c}}$ to reach $T_{\mathrm{BEC}}$. Possible routes to evade these constraints include quantum geometric and hybridization related band structure effects 90, 91, 92, 93, 94, 95, 96, 97, 98 as well as the multiorbital Hund’s interaction effects triggering substantial localized stiffness uncovered here.
Table 1: Characteristics of the BCS, crossover to lattice BEC, and multiorbital strong coupling limits. Behavior of the critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, superconducting stiffness $D_{\mathrm{s}}$, and the ratio of superconducting gap $\Delta$ to the Fermi energy $E_{\mathrm{F}}$ (coupling strength) in the BCS, crossover to lattice BEC 24 and “multiorbital strong coupling” limit. We indicate in the last row where we can find the respective behavior in the interaction plane of the A3C60 model. We do not observe the deep BEC limit for large $U/W$ and small $J/W$, but signatures indicating the onset of the BCS–BEC crossover (c.f. Figs. 4a and 5 as well as Supplementary Note 6).
  BCS limit Crossover to lattice BEC limit Multiorbital strong coupling
$\bm{T_{\mathrm{c}}}$ Increase with pairing interaction, $T_{\mathrm{c}}\propto\Delta(T=0)$ Decrease with pairing interaction, $T_{\mathrm{c}}\propto D_{\mathrm{s}}$ Increase with pairing interaction, $T_{\mathrm{c}}\propto\Delta,\,D_{\mathrm{s}}$
$\bm{\xi_{0}}$ Long ($\xi_{0}\,\gg\,a$) Short ($\xi_{0}\,\gtrsim\,a)$ increasing in (deep) BEC limit Short ($\xi_{0}\sim a)$ constant with pairing interaction
$\bm{D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}}$ Constant with pairing interaction Decreasing with pairing interaction Increasing/constant with pairing interaction
$\bm{\Delta}$/$\bm{E_{\mathrm{F}}}$ Small value ($\Delta/E_{\mathrm{F}}\ll 1$) Large value ($\Delta/E_{\mathrm{F}}\gtrsim 1$) Intermediate value ($\Delta/E_{\mathrm{F}}\sim\mathcal{O}(0.1)$)
Corresponding ($\bm{U}$, $\bm{J}$) region $\sim$ small $U/W$ $\sim$ small $|J|/W$ Onset at $\sim$ large $U/W$ $\sim$ small to intermediate $U/W$ $\sim$ large $|J|/W$, $|J|/U<1/2$
We emphasize that the framework to calculate the coherence length $\xi_{0}$ and the London penetration depth $\lambda_{\mathrm{L}}$ introduced in this work can be implemented in any Green function or density functional based approach to superconductivity without significant increase of the numerical complexity. Thus, our work opens the gate for “in silico” superconducting materials’ optimization targeting not only $T_{\mathrm{c}}$ but also $\xi_{0}$ and $\lambda_{\mathrm{L}}$. On a more fundamental level, availability of $\xi_{0}$ and $\lambda_{\mathrm{L}}$ rather than $T_{\mathrm{c}}$ alone can provide more constraints on possible pairing mechanisms through more rigorous theory-experiment comparisons, particularly in the domain of superconductors with strong electronic correlations.
Methods
Dynamical mean-field theory under the constraint of finite-momentum pairing
We study the multiorbital interacting model $H_{\mathrm{kin}}+H_{\mathrm{int}}$ of A3C60 using Dynamical Mean-Field theory (DMFT) in Nambu space under the constraint of finite-momentum pairing (FMP). In DMFT, the lattice model is mapped onto a single Anderson impurity problem with a self-consistent electronic bath that can be solved numerically exactly. The self-energy becomes purely local $\Sigma_{ij}(i\omega_{n})=\delta_{ij}\Sigma(i\omega_{n})$ capturing all local correlation effects. To this end, we have to solve the following set of self-consistent equations
[…]
in DMFT 51. Here, the local Green function $G_{\mathrm{loc}}$ is obtained from the lattice Green function $G_{\bm{k}}$ (first line) in order to construct the impurity problem by calculating the Weiss field $G_{\mathrm{W}}$ (second line). Solving the impurity problem yields the impurity Green function $G_{\mathrm{imp}}$ from which the self-energy $\Sigma$ is derived (third line). Underlined quantities denote matrices with respect to orbital indices ($\alpha$), $\mathds{1}$ is the unit matrix in orbital space, $\omega_{n}=(2n+1)\pi T$ denote Matsubara frequencies, $h_{\alpha\gamma}(\bm{k})=\sum_{j}t_{\alpha\gamma}(\bm{R}_{j})\mathrm{e}^{i\bm{k}\bm{R}_{j}}$ is the Fourier transform of the hopping matrix in Eq. (8), $\mu$ indicates the chemical potential, and $N_{\bm{k}}$ is the number of $\bm{k}$-points in the momentum mesh. Convergence of the self-consistency problem is reached when the equality $\underline{G}_{\mathrm{loc}}(i\omega_{n})=\underline{G}_{\mathrm{imp}}(i\omega_{n})$ holds.
We can study superconductivity directly in the symmetry-broken phase by extending the formalism to Nambu–Gor’kov space. The Nambu–Gor’kov function under the constraint of FMP (c.f. Eq. (4)) on Matsubara frequencies
$[\underline{\mathcal{G}}_{\bm{q}}(i\omega_{n},\bm{k})]^{-1}$ $=\begin{pmatrix}(i\omega_{n}+\mu)\mathds{1}-\underline{h}(\bm{k}+\frac{\bm{q}}{2})-\underline{\Sigma}^{\mathrm{N}}(i\omega_{n})&-\underline{\Sigma}^{\mathrm{AN}}(i\omega_{n})\\ -\underline{\Sigma}^{\mathrm{AN}}(i\omega_{n})&(i\omega_{n}-\mu)\mathds{1}+\underline{h}(-\bm{k}+\frac{\bm{q}}{2})+[\underline{\Sigma}^{\mathrm{N}}]^{*}(i\omega_{n})\end{pmatrix}$ (10)
then takes the place of the lattice Green function $\underline{G}_{\bm{k}}(i\omega_{n})\mapsto\underline{\mathcal{G}}_{\bm{q}}(i\omega_{n},\bm{k})$ in the self-consistency cycle in Eq. (9). In addition to the normal component $\Sigma^{\mathrm{N}}\equiv\Sigma$, the self-energy gains an anomalous matrix element $\Sigma^{\mathrm{AN}}$ for which the gauge is chosen such that it is real-valued.
We use a $35\times 35\times 35$ $\bm{k}$-mesh and 43200 Matsubara frequencies to set up the lattice Green function in the DMFT loop. In order to solve the local impurity problem, we use a continuous-time quantum Monte Carlo (CT-QMC) solver 99 based on the strong coupling expansion in the hybridization function (CT-HYB) 100. Details on the implementation can be found in Refs. 67, 68. Depending on calculation parameters ($T$, $U$, $J$) and proximity to the superconducting transition, we perform between $2.4\times 10^{6}$ up to $19.2\times 10^{6}$ Monte Carlo sweeps and use a Legendre expansion with 50 up to 80 basis functions. Some calculations very close to the onset of the superconducting phase transition (depending on $T$ or $\bm{q}$) needed more than 200 DMFT iterations until convergence. We average 10 or more converged DMFT iterations and calculate the mean value and standard deviation in order to estimate the uncertainty of the order parameter originating from the statistical noise of the QMC simulation.
Further details on code implementation are given in Supplementary Note 4. It entails a comparison of finite-momentum pairing to the zero-momentum case51 in the Nambu–Gor’kov formalism, an explanation on the readjustment of the chemical potential $\mu$ to fix the electron filling, and a simplification of the Matsubara sum for calculating $\bm{j}_{\bm{q}}$ in Eq. (6).
Acknowledgements.
We thank M. Capone, A. Cavalleri, M. Eckstein, M. Katsnelson, Y. Iwasa, L. Mathey, T. Nomoto, G. Rai, S. Ryee, A. Toschi, and P. Werner for fruitful discussions. N.W., A.I.L., and T.W. acknowledge funding by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) (EXC 2056, Project ID 390715994) and the DFG research unit FOR 5249 (‘QUAST’, Project No. 449872909). Y.N. is grateful to funding via JSPS KAKENHI (No. JP23H04869 and JP21H01041). R.A. acknowledges funding by JSPS KAKENHI (No. JP24H00190) and RIKEN-TRIP initiative (Many-body Electron Systems). A.I.L. is supported by the European Research Council via Synergy Grant 854843 - FASTCORR. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board under the project hhp00056 and provided on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure. Additional calculations were performed on the Physnet cluster at Universität Hamburg. We acknowledge financial support from the Open Access Publication Fund of Universität Hamburg.
Author contributions: N.W. performed the calculations, prepared the figures and wrote the manuscript with input from all authors. Y.N. and R.A. provided the code for the DMFT calculations of the A3C60 model which A.I.L. and N.W. modified to include FMP. N.W. and S.B. analyzed the GL formulation of the FMP constraint. All authors discussed the results and physical implications. T.W., A.I.L., and N.W. conceived the project. T.W. supervised the project.
Competing interests: All authors declare that they have no competing interests.
Data availability: The data that support the findings of this study are available from the corresponding author upon reasonable request.